you by 37 concert tickets for eight dollars each and then solve 37 tickets for 11 each word below shows two ways you can determine your profit this game it’s a solution method which do you prefer explain your reasoning
Ex.1 profit= 37(11)- (37)8
= 407-296
= 37(3)
= $111
Ex.2 profit =37(11)-(37)8
= 407-296
= $111

Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52847)

.

I prefer BOTH.


-------------

Hey, this question is addressed to you: they ask YOU about YOUR opinion.

You re-address this question to me.

My opinion can be opposite to your, and, reversely, your opinion can be opposite to my.

So, I simply do not understand the GAME you play in . . .



Answer by MathTherapy(10555)


you by 37 concert tickets for eight dollars each and then solve 37 tickets for 11 each word below shows two ways you can determine your profit this game it’s a solution method which do you prefer explain your reasoning
Ex.1 profit= 37(11)- (37)8
= 407-296
= 37(3)
= $111
Ex.2 profit =37(11)-(37)8
= 407-296
= $111
The former is better, because it's easier to calculate, especially when in an exam that is DEVOID of a calculator, and time is of the essence.
The former involves less calculations, as you can see (37 * 3), as opposed to the latter which involves calculating 37 * 11, and 37 * 8 first,
and then subtracting them. Most would not want to do all this work, as it's time-consuming, and involves too much, to those folks!